what is the answer to this problem? 25+6-15+25-10+14-21
how do I solve this problem. Please set it up for me thank you.
A $100.00 check that is cashed at a rate of $7.00, at what rate did you cash it? (in percentage)

Answer by Alan3354(69443)

what is the answer to this problem? 25+6-15+25-10+14-21
how do I solve this problem. Please set it up for me thank you.
I assume you want the sum of those numbers.
One way to do it is to add all the pos and all the neg numbers first.
= 25+6+25+14 + (-15-10-21)
= 70 + (-46)
= 70 - 46
= 24
-------
A $100.00 check that is cashed at a rate of $7.00, at what rate did you cash it? (in percentage)
I don't know what that means. $7 is not a rate, it's an amount.
$7 is 7% of $100, if that's your question.